a jar contains 7 green, 19 black and 13 pik marbles. A marble is drawn at random. What is the probability of it being a green or black?
Answer by ewatrrr(24785)


Hi
a jar contains 7 green, 19 black and 13 pink marbles.
39 marbles in ALL
ONE marble is drawn at random.
P(being a green 'or' black) = 7/39 + 19/39 = 26/39